the area of a swimming pool is 200 square feet. the lengthof the pool is 10 feet longer than the width.what is the width of the swimming pool?

Answer by nerdybill(7384)

the area of a swimming pool is 200 square feet. the lengthof the pool is 10 feet longer than the width.what is the width of the swimming pool?
.
Let w = width of pool
then
w+10 = length of pool
.
w(w+10) = 200
w^2 + 10w = 200
w^2 + 10w - 200 = 0
(w+20)(w-10) = 0
w = {-20, 10}
you can throw out the negative solution (extraneous) leaving:
w = 10 feet